A plane can travel 640 miles against the wind in the same time it can travel 800 miles with the wind. If the plane can fly 180 mph in calm air, what is the speed of the wind?
Answer by checkley77(12844)

D=RT OR T=D/R
T=640/(180-W)
T=800/(180+W)
640/(180-W)=800/(180+W)
640(180+W)=800(180-W)
115,200+640W=144,000-800W
640W+800W=144,000-115,200
1,440W=28,800
W=28,800/1,440
W=20 MPH IS THE WIND SPEED.
PROOF:
640/(180-20)=800/(180+20)
640/160=800/200
4=4